a) Calcuate the pH of a solution with [H3O+] = 6.54 x 10-3 M

b) Calcuate the pOH of a solution with [H3O+] = 8.5 x 10-9 M

c) A solution has a pH = 2.420, what is the [OH-]?

d) A solution has a pOH = 1.0, what is the [OH-]?

e) Calcuate the pH of a solution with [OH-] = 5.9 x 10-6 M
Answer #1

sylewejgxə 14=? +H+) = (20+) = 65481044 Pr= -day (H+] = -log (6.5 4X107 hotea - I loy [854) 4 dog lo?? the dog (6.54) +24910Syаеше хә ☺ fra 2.420 totta? (+= 10-4H 10 24 (pt) & (op] = 10-14 204) = 107th / LA47 = 10-14 10-242 LOH] = 10-1412-14 Lox) =☺ [04] = 5.9x10-6M pOH = 1:0 pore= -log[ou ] 1 = -log[on-] [on] antilag (-1) O.IM pOH = -logion ] pot = -log (5.9x10 m] [dosi